# Unified Triplet-Level Hallucination Evaluation for Large Vision-Language Models

### v3

@@ 1 Introduction
With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination. However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning (Gou et al., 2025a). As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image. Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
2) Hallucination discrimination: To evaluate how severe LVLMs hallucinate objects and their relationships within given images, prior works generally use either self-discrimination methods (e.g., Yes/No questions) (Li et al., 2023e; Wang et al., 2023b; Guan et al., 2024; Wu et al., 2024b) or template-driven discrimination approaches (e.g., “What is the relation with A and B?”) such as Reefknot (Zheng et al., 2024). However, such methods inherently constrain LVLMs to generate short answers like “Yes/No” or “A has {} relation to B”. Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024). For instance, Li et al. (2023e) have shown that InstructBLIP (Dai et al., 2023) tends to produce shorter outputs compared to other LVLMs, thus inflating its performance in answering the above type of questions and leads to hallucination evaluation bias. Moreover, these benchmarks require transforming general vision-language tasks into specific formats like "Yes/No", limiting their applicability. Therefore, we raise the following research question: Can we develop a unified and unbiased evaluation framework capable of evaluating various types of hallucinations in LVLM responses across diverse tasks?
@@ 5.2 Main Result
Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024)
@@ Results.
Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al. (2023a) and MiniGPT-4 Zhu et al. (2023) with baseline comparison. The best results under each column are boldfaced.
@@ Appendix C Human Annotation Guideline
Appendix D More Discussion
Future works.
Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2023c; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning.

### v4

@@ 1 Introduction
With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination. However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting reasoning over these objects(Gou et al., 2025a; Yu et al., 2025). As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image. Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
2) Hallucination discrimination: To evaluate how severe LVLMs hallucinate objects and their relationships within given images, prior works generally use either self-discrimination methods (e.g., Yes/No questions) (Li et al., 2023e; Wang et al., 2023b; Guan et al., 2024; Wu et al., 2024b) or template-driven discrimination approaches (e.g., “What is the relation with A and B?”) such as Reefknot (Zheng et al., 2024). However, such methods inherently constrain LVLMs to generate short answers like “Yes/No” or “A has {} relation to B”. Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024b). For instance, Li et al. (2023e) have shown that InstructBLIP (Dai et al., 2023) tends to produce shorter outputs compared to other LVLMs, thus inflating its performance in answering the above type of questions and leads to hallucination evaluation bias. Moreover, these benchmarks require transforming general vision-language tasks into specific formats like "Yes/No", limiting their applicability. Therefore, we raise the following research question: Can we develop a unified and unbiased evaluation framework capable of evaluating various types of hallucinations in LVLM responses across diverse tasks?
@@ 5.2 Main Result
Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024). Check Appendix D for evaluation results on more LVLMs.
@@ Results.
Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al. (2023a) and MiniGPT-4 Zhu et al. (2023) with baseline comparison. The best results under each column are boldfaced. Check more baseline comparison results in Appendix D.
@@ Appendix C Human Annotation Guideline
Appendix D Additional Results
In this section, we present additional evaluation results to supplement those in Table 2 and Table 5, enabling a more comprehensive evaluation. Due to the deprecation of gpt-4-1106-preview—previously employed in LLM Judge (GPT-4 Triplet)—by OpenAI’s API, and our limited budget for querying proprietary models, we adopt LLM Judge (Llama-3.3 Triplet) for the following experiments. Its effectiveness has been demonstrated in Table 3.
D.1 Evaluating more recent LVLMs on Tri-HE
To ensure a more comprehensive evaluation, we further evaluate InternVL2_5-8B Chen et al. (2024e) and Qwen2.5-VL-7B-Instruct Team (2025) on Tri-HE. Since we use the LLM Judge (LLaMA-3.3 Triplet) in this setting—which differs from the judge used in Table 2—we also report the results of InternLM-X2 and LLaMA-3.2 under the same judge for comparison.
The results are presented in Table 9. As shown, all recent models still exhibit a certain degree of both object and relation hallucinations, demonstrating the effectiveness of Tri-HE in identifying hallucination issues in LVLMs. Notably, InternVL2_5-8B achieves the lowest hallucination rates across all metrics, suggesting that its superior pre-training data quality and the use of Random JPEG Compression as a data augmentation strategy are effective in mitigating hallucinations in model responses.
Method LLM Judge
Overall Object Relation
$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
LLaMA-3.2 38.98 37.37 23.25 21.13 13.51 13.30
Qwen2.5-VL-7B-Instruct 38.19 35.92 26.16 24.09 11.83 11.38
InternLM-X2 31.75 29.93 19.29 18.09 12.46 11.84
InternVL2_5-8B 29.12 26.99 17.16 14.95 12.22 11.13
Table 9: Comparison on hallucination rates among recent LVLMs on Tri-HE with LLM Judge (Llama-3.3 Triplet).
D.2 More Hallucination Mitigation Methods
In this section, we include additional recent training-free methods for hallucination reduction to enable a more comprehensive comparison. Specifically, we compare our mitigation framework with a decoding-based approach, VCD (Leng et al., 2024). As shown in Table 10, although VCD is designed to address object hallucination, it exhibits more severe hallucination issues when evaluated on our reasoning-grounded question set compared to our proposed mitigation method.
In addition, reinforcement learning (RL) approaches have received growing attention for addressing hallucination. To enable a more comprehensive evaluation, we assess a RL-trained method, OPA-DPO (Yang et al., 2025), using our evaluation framework. The results in Table 10 demonstrate its effectiveness. It is worth noting that, since OPA-DPO requires additional training, its performance advantage over our method is expected. Nonetheless, it can potentially be integrated with our proposed approach as a foundation for further RL fine-tuning.
Model Mitigation Type $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
LLaVA-1.5 w/o Mitigation - 21.73 21.07
VCD training-free 28.89 30.64
Ours training-free 19.53 20.98
OPA-DPO RL-trained 14.44 16.82
Table 10: Additional baseline comparisons on hallucination mitigation with Llama3.3 Judge. The best results under each column are boldfaced.
Appendix E Future works
Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2024a; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning. Furthermore, additional types of hallucination can be incorporated into our triplet-level evaluation framework, such as temporal changes of an object over time, represented in the form of (old, change, new) triplets. We plan to explore these extensions in future work.